PointsU1U2

5. Imagine there is an economy with pure exchange with two individuals (1 and 2). A, B and C are three points that belong to the utility possibility frontier (UPF) of the two individuals.

a) According to the information of the previous table, is the following statement true or false: "The point D (2,5) belong to the UPF"

b) Please explain why it is true or false.

c) If the Social Welfare Function (SWF) of this economy is W= U1+ U2. Which of the points A, B, C or D maximizes the SWF?

d) Write the letter of the point that maximizes the SWF.
Answer #1

A) No, it will lie inside the UPF since the point A (2,7) which is clearly greater than D(2.5) lies on the UPF

B) False (explained above)

C) W(A) = 2+7 = 9

W(B) = 5+5 = 10

W(C) = 7+2 = 9

W(D) = 2+5 = 7

Thus, allocation A maximizes SWF

D) A